OpenStudy (anonymous):

solve: a^2 · a^3 · a^-4

OpenStudy (anonymous):

Nothing to solve... maybe they mean simplify? So when you multiply a variable with different exponents, what do you do the exponents... You add them right? For example: \[\large x^5\cdot x^6 = x^{5+6} = x^{11}\]

OpenStudy (anonymous):

I cant see the example sorry, but would the answer be a^-1?

OpenStudy (anonymous):

i mean a^1

OpenStudy (anonymous):

Yeah, but remember that \(a^1\) is just \(a\).

OpenStudy (anonymous):

yes this answer is right..

OpenStudy (anonymous):

(a^2+3) (a^-4) a^5 . a^-4 a^5-4 = a^1

OpenStudy (anonymous):

ty